OpenStudy (kaynick):

The points (−3,7) and (1,r) lie on a line with slope −1. Find the missing coordinate r.

OpenStudy (anonymous):

The gradient of a line is found by x2 - 1/y2 - y1 1-- 3/r- 7 = -1 4/r-7 = -1 4= -r + 7 r= 3

OpenStudy (kaynick):

thank you :)

OpenStudy (anonymous):

Welcome